OpenStudy (anonymous):

Is \(0i\in \mathbb C \setminus \mathbb R\)?

OpenStudy (anonymous):

If you edit it it doesn't render any more, lol.

OpenStudy (anonymous):

Is \( \Large 0i\in \mathbb C \setminus \mathbb R\)?

OpenStudy (kainui):

This means is zero times sqrt(-1) an element of the complex numbers and not an element of the reals? Jeeze, I honestly don't know haha. How are the complex numbers defined?

OpenStudy (anonymous):

Complex numbers are defined as \(a+bi\) where \(a,b\in \mathbb R\) and \(i=\sqrt{-1}\)

OpenStudy (kainui):

Is 0=0i? Since usually I see it as a+bi, I'm going to have to say that it is an element, since b is a real number times i, so really when we see 0 it is assumed to be 0+0i?

OpenStudy (kainui):

Yeah not very proof like of me. So maybe: a is an element of R a and b are an elements of C then C\R is only b. Since 0 is a possible b, 0i is contained in C\R

OpenStudy (kainui):

I'm in Real Analysis 1 right now, so give me a break, but I feel like that's pretty much what you're looking for, it needs more formalization going on there though haha.

OpenStudy (kainui):

This is kind of like saying the zero vector is contained in every coordinate system, but if you remove one of the axes, you're only removing the 0 in that direction. Similarly, you don't gain an extra point if you add another dimension either. Sort of weird but makes sense if you think about it.

OpenStudy (zarkon):

\[0i=0\in\mathbb{R}\]